# Federated Unlearning: a Perspective of Stability and Fairness

### v1

@@ 5.1 Trade-off between FU Verification and Stability
$C_{1}=$ $\left(1+2\frac{\beta_{\mathcal{S}}^{2}-1}{T}\right)\left(\Delta F_{-\mathcal{J}}\left(\boldsymbol{w}^{*},\boldsymbol{w}^{r*}\right)\right.$ (5)
$\left.+\Delta F_{-\mathcal{J}}\left(\boldsymbol{\boldsymbol{w}}^{o},\boldsymbol{w}^{*}\right)\right)+\frac{1}{2LT}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$
@@
The effectiveness of FU, as measured by $V(\boldsymbol{{w}}^{u})$, is hindered by its lower bound $C_{1}$ in Equation 5 with several factors:
@@
Let Assumptions 5.1-5.5 hold, and given an original trained model $\boldsymbol{\boldsymbol{w}}^{o}$ undergoing unlearning. Consider the learning rate $\eta=\frac{1}{T\sqrt{\mu}}\sqrt{\frac{\beta_{\mathcal{S}}-1}{\min\left\{\mu(\beta_{\mathcal{S}}-1),L(\beta_{\mathcal{S}}-1)\right\}}}$, the sum of the unlearning verification metric and the global stability metric $V(\boldsymbol{{w}}^{u})+S(\boldsymbol{{w}}^{u})$ is bounded below by a constant $C_{s}$, i.e.,
$C_{s}=\frac{P_{\mathcal{J}}}{\sqrt{2}\mu}\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\|^{2}+\delta+C_{1},$ (7)
where $C_{1}$ is defined in Lemma 5.8.
Theorem 5.10 illustrates a fundamental trade-off in FU: effectively unlearning clients ($V$) while maintaining the stability of the global model’s performance ($S$). This trade-off is determined by the divergence between the optimal model $\boldsymbol{w}^{*}$ and the optimal unlearned model $\boldsymbol{w}^{r*}$. Specifically, achieving effective FU and stability is not feasible under the substantial divergence between $\boldsymbol{w}^{*}$ and $\boldsymbol{w}^{r*}$.
@@ 5.2 Trade-off between FU Verification and Fairness
Within FU, the sum of the unlearning verification metric and the local fairness metric $2V(\boldsymbol{{w}}^{u})+Q(\boldsymbol{{w}}^{u})$ is bounded below by a constant $C_{q}$:
@@ 6 Optimizing FU under Trade-offs
6.1 FU for Balancing Global Stability
In FU, maintaining global stability is crucial for ensuring the overall performance and reliability of the federated system throughout the unlearning process. However, as explored in Section 5.1, a trade-off exists between FU verification and global stability. To manage this trade-off, we propose an FU mechanism utilizing a penalty-based approach and gradient correction techniques. We also theoretically demonstrate the convergence of our method.
FU Mechanism Design: To balance stability during FU, we formulate the optimization problem for unlearning as:
@@ 6.1 FU for Balancing Global Stability
By adjusting $\lambda$, we can manage the trade-off between these two objectives, allowing for a flexible approach to specific requirements of the federated system.
By the definition of stability metric $S(\boldsymbol{w})$, we have: $S(\boldsymbol{w})=\mathbb{E}\left[F(\boldsymbol{w})\right]-F(\boldsymbol{w}^{*})=\mathbb{E}\left[F(\boldsymbol{{w}}^{u})\right]-F(\boldsymbol{\boldsymbol{w}}^{o})+\delta,$ where $\delta=F(\boldsymbol{\boldsymbol{w}}^{o})-F(\boldsymbol{w}^{*})$. Consequently, solving P2: $\min_{\boldsymbol{w}}F_{-\mathcal{J}}(\boldsymbol{w})+\lambda\left(F(\boldsymbol{w})-F(\boldsymbol{\boldsymbol{w}}^{o})\right)$ optimizes P1. However, in P2, optimizing the global objective $F(\boldsymbol{w})$ among all clients is untraceable in FU as it cannot involve unlearned client $j\in\mathcal{J}$ in unlearning process. To address this, we consider the approximate problem P3 to P2:
[…]
where $\tilde{h}(\boldsymbol{w})=(1-P_{\mathcal{J}})F_{1}+P_{\mathcal{J}}F_{2}$, $F_{1}:=\lambda F_{-\mathcal{J}}(\boldsymbol{w})$, and $F_{2}:=\lambda\left(\langle\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o}),\boldsymbol{w}-\boldsymbol{\boldsymbol{w}}^{o}\rangle+\frac{L}{2}\left\|\boldsymbol{w}-\boldsymbol{\boldsymbol{w}}^{o}\right\|^{2}\right)$.
To address P3, we propose an FU mechanism that operates two steps during each unlearning round $t$:
@@
The gradients of local objectives are bounded, i.e., $\|\nabla f_{i}(\boldsymbol{w})\|\leq G$ for all $i$. This implies the gradient of global objective $F_{-\mathcal{J}}$ for remaining clients is also bounded: $\|\nabla F_{-\mathcal{J}}(\boldsymbol{w})\|\leq G$.
@@
The heterogeneity between the unlearned clients $\mathcal{J}$ and the remaining clients is quantified as:
${\left\|\mathbb{E}\left[\boldsymbol{\hat{g}}_{\mathcal{J}}^{t}(\boldsymbol{w})\right]-\nabla F_{-\mathcal{J}}(\boldsymbol{w})\right\|^{2}}\leq\zeta^{\prime 2}+\beta^{\prime 2}\left\|\nabla F_{-\mathcal{J}}(\boldsymbol{w})\right\|^{2},$
where $\boldsymbol{\hat{g}}_{\mathcal{J}}^{t}(\boldsymbol{w})=\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})+L(\boldsymbol{w}-\boldsymbol{\boldsymbol{w}}^{o})$, $\zeta^{{}^{\prime}2}$ and $\beta^{{}^{\prime}2}$ indicate heterogeneity between unlearned and remaining clients.
@@
Under 6.2, the expected norm of the gradient correction at unlearning round $t$ is bounded:
$\mathbb{E}\left[\left\|\boldsymbol{g_{c}}({\boldsymbol{\bar{w}}}^{(t,E)})\right\|^{2}\right]\leq\phi\left(\zeta^{\prime 2}+(\beta^{\prime 2}+1)\left\|\nabla F_{-\mathcal{J}}({\boldsymbol{\bar{w}}}^{t,E})\right\|^{2}\right),$
where $\phi=\lambda^{2}P_{\mathcal{J}}^{2}(1+\cos^{2}\theta)$, $\cos^{2}\theta$ represents the similarity in objectives between remaining and unlearned clients.
@@
$\mathbb{E}\left[H\left(\bar{\mathbf{w}}^{(t+1,0)}\right)\right]-H\left(\mathbf{w}^{*}\right)\leq\frac{L}{2}\left(\frac{v}{\gamma+t}+\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{*}\right\|^{2}\right),$
where $v=\max\left\{\frac{\beta^{2}B}{\beta\mu-4},(\gamma+1)\left\|\boldsymbol{\boldsymbol{w}}^{o}-\boldsymbol{w}^{r*}\right\|^{2}\right\}$.
Our approach introduces additional complexity in the convergence analysis compared to that of Li et al. (2020b, Theorem 1) due to incorporating a gradient correction in FU, as detailed in Section E.1 This complexity is reflected in $B$ with additional components: $2\phi\left(\frac{\eta_{g}}{\eta_{l}}\right)^{2}((\beta^{\prime 2}+1)G^{2}+\zeta^{\prime 2})$. It highlights two insights:
@@
In the special case of homogeneity between remaining and unlearned clients, where the original model and the optimal unlearned model are ideally aligned ( $\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{*}\right\|^{2}=0$, $\beta^{\prime 2}=\zeta^{\prime 2}=0$, $\cos\theta=1$, and $P_{\mathcal{J}}=0.5$), the additional term in $B$ reduces to $\lambda^{2}\left(\frac{\eta_{l}}{\eta_{g}}\right)^{2}G^{2}$. In this scenario, handling stability in FU is straightforward for the tight bound. Conversely, our mechanism reduces the convergence bound in heterogeneous settings, characterized by orthogonal gradient correction to remaining clients’ gradients ($\cos\theta=0$). This indicates we effectively adapt to this heterogeneity.
@@
Under Assumptions 5.1-5.4, and Assumptinos 6.1-6.5, taking $\eta_{l}=\eta_{g}=\frac{2}{LT}$, and $T\geq\max\left\{2\beta_{\mathcal{S}}^{2}+2,\frac{1+\Delta}{4L},\frac{1}{2}(\beta^{\prime\prime 2}+1),\frac{1+\Delta^{\prime}}{L}\right\}$, where $\Delta=\sqrt{\max\left\{0,1-16L(\beta_{\mathcal{S}}^{2}+1)\right\}}$ and $\Delta^{\prime}=\sqrt{\max\left\{0,1-L(\beta^{\prime\prime 2}+1)\right\}}$. Then, the verification metric $V$ is bounded as follows:
$V=\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{{w}}^{u})-F_{-\mathcal{J}}(\boldsymbol{w}^{*})\right]\leq\chi_{1}+\chi_{2}$, where
$\chi_{1}$ $=\frac{1}{2}\left(1-\frac{1}{2LT}+\frac{\beta_{\mathcal{S}}^{2}+1}{LT^{2}}\right)^{T}D+\frac{(\sigma^{2}+\zeta_{\mathcal{S}}^{2})}{2LT},$
$\chi_{2}$ $=\frac{1}{2}\left(1-\frac{1}{LT}+\frac{\beta^{{}^{\prime\prime}2}+1}{LT^{2}}\right)^{T}D+\frac{\zeta^{\prime\prime 2}}{2LT}.$
Here, $D=F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*}).$
@@ 6.2 FU for Balancing Local Fairness
P4: $\min_{\boldsymbol{w}}F_{-\mathcal{J}}(\boldsymbol{w})$
[…]
To solve this problem, we adapt the saddle point optimizations as in Agarwal et al. (2018); Hu et al. (2022), using a Lagrangian multiplier $\lambda_{i}$ for each constraint:
[…]
where $\boldsymbol{r}(\boldsymbol{\boldsymbol{w}}^{o})=\left[\Delta f_{i}-\epsilon\right]_{i\notin\mathcal{J}}$. The detailed algorithm for solving this problem is provided in the Appendix F.
@@
This lemma can be derived from Hu et al. (2022, Theorem 1). It indicates the required unlearning rounds to reach a $\nu$-approximate saddle point of P5. It highlights that increased data heterogeneity among remaining clients and stringent fairness constraints require more unlearning rounds $T$ to balance the trade-off.
@@
Given $\epsilon=\frac{1}{\Lambda}(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}^{*})$ and assuming the existence of $\nu$-approximate saddle points of the trade-off fairness problem, then the unlearning verification metric $V\leq 2\nu$ and $\max_{i\notin\mathcal{J}}\Delta f_{i}\leq\epsilon$.
@@
    Data Hetegeneity: When the original FL model and optimal FU model are homogeneous, then $\epsilon=F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{*})$, there is no fairness loss from unlearning but from data heterogeneity among remaining clients (as discussed in Theorem 5.11). However, with higher heterogeneity and lacking a focus on balancing fairness (characterized by a negligible $\Lambda$ and a small $\nu$), FU compromises fairness $\epsilon$ to reduce $V$.
@@
Figure 4: Unlearning Convergence and Handling Stability: Increasing stability penalty $\lambda$ enhances global performance of FU.
7 Experiments

### v4

@@ 5.1 Trade-off between FU Verification and Stability
$C_{1}=\left(1+{\frac{\beta_{\mathcal{S}}^{2}-1}{T}}\right)\left(\Delta F_{-\mathcal{J}}\left(\boldsymbol{w}^{*},\boldsymbol{w}^{r*}\right)\right.\left.\,+\Delta F_{-\mathcal{J}}\left(\boldsymbol{\boldsymbol{w}}^{o},\boldsymbol{w}^{*}\right)\right)+\frac{1}{2LT}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right)\,,$
@@
The effectiveness of FU, as measured by $V(\boldsymbol{{w}}^{u})$, is hindered by its lower bound $C_{1}$ in Lemma 5.6 with several factors:
@@
Let Assumptions 5.1-5.5 hold, and given an original trained model $\boldsymbol{\boldsymbol{w}}^{o}$ with FU mechanism $\mathcal{M}$. Consider the learning rate $\eta=\frac{1}{T\sqrt{\mu}}\sqrt{\frac{\beta_{\mathcal{S}}-1}{\min\left\{\mu(\beta_{\mathcal{S}}-1),L(\beta_{\mathcal{S}}-1)\right\}}}$, the sum of the FU verification metric and the global stability metric $V(\boldsymbol{{w}}^{u})+S(\boldsymbol{{w}}^{u})$ is bounded below by:
$V(\boldsymbol{{w}}^{u})+S(\boldsymbol{{w}}^{u})\geq C_{s}=\frac{P_{\mathcal{J}}}{\sqrt{2}\mu}\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\|^{2}+\delta+C_{1}\,,$ (6)
where $C_{1}$ is defined in Lemma 5.8 and $C_{s}$ is constant.
Theorem 5.10 illustrates a fundamental trade-off in FU: effectively unlearning clients ($V$) while maintaining the stability of the global model’s performance ($S$). This trade-off is determined by the divergence between $\boldsymbol{w}^{*}$ and $\boldsymbol{w}^{r*}$. Specifically, achieving effective FU and stability is not feasible under the substantial divergence between $\boldsymbol{w}^{*}$ and $\boldsymbol{w}^{r*}$.
@@ 5.2 Trade-off between FU Verification and Fairness
Within FU, the sum of the unlearning verification metric and the local fairness metric is bounded below by a constant $C_{q}$:
@@ 6 Optimizing FU under Trade-offs
Algorithm 2 FU Mechanism Design under
Global Stability Constraint.
1:  Input: $\boldsymbol{\boldsymbol{w}}^{o}$ (original model), $\mathcal{J}$ (unlearn clients), $T$ (unlearn rounds), $E$ (local epochs), $\eta_{l},\eta_{g}$ (learning rates), $\lambda$ (stability penalty)
2:  Initialize $\boldsymbol{w}_{0}\leftarrow\boldsymbol{\boldsymbol{w}}^{o}$
3:  for $t=0$ to $T-1$ do
4:     $\mathcal{S}\leftarrow$ subset of remaining clients $\mathcal{N}\setminus\mathcal{J}$
5:     for client $i\in\mathcal{S}$ in parallel
6:     $\quad\boldsymbol{w}^{(t,E)}_{i}\leftarrow\text{LocalTrain}(\boldsymbol{w}_{t},\mathcal{D}_{i},E,\eta_{l})$
6:      // 1. Federated Aggregation.
7:     ${\boldsymbol{\bar{w}}}^{(t,E)}\leftarrow\sum_{i\in\mathcal{S}}\alpha_{i}\boldsymbol{w}^{(t,E)}_{i}$
7:      // 2. Gradient Correction.
8:     $\boldsymbol{h}^{t}\leftarrow\lambda(1-P_{\mathcal{J}})\boldsymbol{g}_{\mathcal{S}}^{t}+\lambda P_{\mathcal{J}}\boldsymbol{\hat{g}}_{\mathcal{J}}^{t}$
9:     $\boldsymbol{g}_{c}^{t}\leftarrow\boldsymbol{h}^{t}-\text{Proj}_{\boldsymbol{g}_{\mathcal{S}}^{t}}\boldsymbol{h}^{t}$
10:     $\boldsymbol{w}_{t+1}={\boldsymbol{\bar{w}}}^{(t+1,0)}\leftarrow{\boldsymbol{\bar{w}}}^{(t,E)}-\eta_{g}\boldsymbol{g}_{c}^{t}$
11:  end for
12:  return $\boldsymbol{{w}}^{u}\leftarrow\boldsymbol{w}_{T}$ // Unlearned model
Algorithm 3 FU Mechanism Design under
Local Fairness Constraint.
1:  Input: $\boldsymbol{\boldsymbol{w}}^{o}$, $\mathcal{J}$, $T$, $E$, $\eta$, $\epsilon$ (fairness threshold), $\Lambda$ (fairness penalty)
2:  Initialize $\boldsymbol{\mu}=\boldsymbol{0}$, $\boldsymbol{w}_{0}\leftarrow\boldsymbol{\boldsymbol{w}}^{o}$
3:  for $t=0$ to $T-1$ do
4:     $\mathcal{S}\leftarrow$ subset of remaining clients $\mathcal{N}\setminus\mathcal{J}$
5:     for client $i\in\mathcal{S}$ in parallel do
6:        $r_{i}\leftarrow f_{i}(\boldsymbol{w}_{t})-f_{i}(\boldsymbol{\boldsymbol{w}}^{o})$
7:        for each $\tau=1,\cdots,E:$
8:        $\quad\boldsymbol{w}_{i}^{(t,\tau+1)}\!\leftarrow\!\boldsymbol{w}_{i}^{(t,\tau)}\!-\!\eta\nabla f_{i}(\boldsymbol{w}_{i}^{(t,\tau)})-{\mu_{i}\eta}(\nabla f_{i}(\boldsymbol{w}_{i}^{(t,\tau)})-f_{i}(\boldsymbol{\boldsymbol{w}}^{o}))\,$
9:     end for
10:     $\boldsymbol{w}_{t}={\boldsymbol{\bar{w}}}^{(t+1,0)}\leftarrow\sum_{i\in\mathcal{S}}\alpha_{i}\boldsymbol{w}_{i}^{(t,E)}.$
11:     If $\max_{i}{r}_{i}\leq\epsilon$: return $\boldsymbol{{w}}^{u}\leftarrow\boldsymbol{w}_{t}$
12:     Update $\mu_{i}=\Lambda\frac{\exp\left(r_{i}\right)}{1+\sum_{i\notin\mathcal{J}}\exp\left(r_{i}\right)}$.
13:  end for
14:  return $\boldsymbol{{w}}^{u}\leftarrow\boldsymbol{w}_{T}$
6.1 Balancing FU Verification and Global Stability
In FU, maintaining global stability is crucial for ensuring the overall performance and reliability of the federated system throughout the unlearning process. However, as explored in Section 5.1, a trade-off exists between FU verification and global stability. To manage this trade-off, we propose an FU mechanism utilizing a penalty-based approach and gradient correction techniques (Algorithm 2). We also theoretically demonstrate the convergence of our method, with further experimental validation in Section 7.
FU Mechanism Design: To balance stability during FU, we formulate the optimization problem for unlearning as P1: $\min_{\boldsymbol{w}}V(\boldsymbol{w})+\lambda S(\boldsymbol{w})$. By adjusting $\lambda$, we can manage the trade-off between these two objectives, allowing for a flexible approach to specific stability requirements of FU.
By the definition of stability metric $S(\boldsymbol{w})$, we have: $S(\boldsymbol{w})=\mathbb{E}\left[F(\boldsymbol{w})\right]-F(\boldsymbol{w}^{*})=\mathbb{E}\left[F(\boldsymbol{w})\right]-F(\boldsymbol{\boldsymbol{w}}^{o})+\delta,$ where $\delta=F(\boldsymbol{\boldsymbol{w}}^{o})-F(\boldsymbol{w}^{*})$. Consequently, solving P2: $\min_{\boldsymbol{w}}F_{-\mathcal{J}}(\boldsymbol{w})+\lambda\left(F(\boldsymbol{w})-F(\boldsymbol{\boldsymbol{w}}^{o})\right)$ optimizes P1. However, in P2, optimizing the global objective $F(\boldsymbol{w})$ among all clients is untraceable in FU as it cannot involve unlearned client $j\in\mathcal{J}$ in unlearning process.
@@ 6.1 Balancing FU Verification and Global Stability
To address P3, we propose an FU mechanism operating two steps during each unlearning round $t$:
@@
The gradients of local objectives are bounded, i.e., $\|\nabla f_{i}(\boldsymbol{w})\|\leq G$ for all $i$.
@@
The heterogeneity between the unlearned clients $\mathcal{J}$ and the remaining clients ${\left\|\mathbb{E}\left[\boldsymbol{\hat{g}}_{\mathcal{J}}^{t}(\boldsymbol{w})\right]-\nabla F_{-\mathcal{J}}(\boldsymbol{w})\right\|^{2}}\leq\zeta^{\prime 2}+\beta^{\prime 2}\left\|\nabla F_{-\mathcal{J}}(\boldsymbol{w})\right\|^{2}$, $\zeta^{{}^{\prime}2}$ and $\beta^{{}^{\prime}2}$ indicate heterogeneity between unlearned and remaining clients.
@@
Under 6.2, the expected norm of the gradient correction at unlearning round $t$ is bounded:$\mathbb{E}\left[\left\|\boldsymbol{g_{c}}({\boldsymbol{\bar{w}}}^{(t,E)})\right\|^{2}\right]\leq\phi\left(\zeta^{\prime 2}+(\beta^{\prime 2}+1)\left\|\nabla F_{-\mathcal{J}}({\boldsymbol{\bar{w}}}^{t,E})\right\|^{2}\right)$, where $\phi=\lambda^{2}P_{\mathcal{J}}^{2}(1+\cos^{2}\theta)$, $\cos^{2}\theta$ represents the similarity in objectives between remaining and unlearned clients.
@@
$\mathbb{E}\left[H\!\left(\bar{\mathbf{w}}^{(t+1,0)}\right)\right]\!\!-\!H^{*}\leq L\frac{v}{\gamma+t}+\frac{1}{2L}P_{\mathcal{J}}^{2}\lambda^{2}G^{2}+\frac{L}{2}\left(\left\|\boldsymbol{w}^{r*}-\boldsymbol{\boldsymbol{w}}^{o}\right\|^{2}+\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}\right)\,,$ (8)
Where $\boldsymbol{w}^{H*}=\text{argmin}_{\boldsymbol{w}}H(\boldsymbol{w})$, and $v\!=\!\max\left\{\frac{\beta^{2}B}{\beta\mu-4},(\gamma+1)\left\|\boldsymbol{w}^{r*}-\boldsymbol{\boldsymbol{w}}^{o}\right\|^{2}\right\}$.
Our approach introduces additional complexity in the convergence analysis compared to that of Li et al. [29, Theorem 1] due to incorporating a gradient correction in FU, as detailed in Section E.1 This complexity is reflected in $B$ with additional components: $2\phi\left(\frac{\eta_{g}}{\eta_{l}}\right)^{2}((\beta^{\prime 2}+1)G^{2}+\zeta^{\prime 2})$. It highlights two insights:
@@
In the special case of homogeneity between remaining and unlearned clients, where the original model and the optimal unlearned model are ideally aligned ( $\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{*}\right\|^{2}=0$, $\beta^{\prime 2}=\zeta^{\prime 2}=0$, $\cos\theta=1$, and $P_{\mathcal{J}}=0.5$), the additional term in $B$ reduces to $\lambda^{2}\left({\eta_{l}}/{\eta_{g}}\right)^{2}G^{2}$, and $\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}=0$ in Equation 8 since $\boldsymbol{w}^{r*}$ is the feasible solution of P3. In this scenario, handling stability in FU is straightforward for the tight bound. Conversely, our mechanism reduces the convergence bound in heterogeneous settings, characterized by orthogonal gradient correction to remaining clients’ gradients ($\cos\theta=0$). This indicates we effectively adapt to this heterogeneity.
@@
Under Assumptions 5.1-5.4, and Assumptinos 6.1-6.5, taking $\eta_{l}=\eta_{g}=\frac{2}{LT}$, and $T\geq\max\left\{2\beta_{\mathcal{S}}^{2}+2,\frac{1+\Delta}{4L},\frac{1}{2}(\beta^{\prime\prime 2}+1),\frac{1+\Delta^{\prime}}{L}\right\}$, where $\Delta=\sqrt{\max\left\{0,1-16L(\beta_{\mathcal{S}}^{2}+1)\right\}}$ and $\Delta^{\prime}=\sqrt{\max\left\{0,1-L(\beta^{\prime\prime 2}+1)\right\}}$.
Then, the FU verification guarantees $V=\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{{w}}^{u})-F_{-\mathcal{J}}(\boldsymbol{w}^{*})\right]\leq\chi_{1}+\chi_{2}$, where $\chi_{1}=\frac{1}{2}\left(1-\frac{1}{2LT}+\frac{\beta_{\mathcal{S}}^{2}+1}{LT^{2}}\right)^{T}D+\frac{(\sigma^{2}+\zeta_{\mathcal{S}}^{2})}{2LT},\quad\chi_{2}=\frac{1}{2}\left(1-\frac{1}{LT}+\frac{\beta^{{}^{\prime\prime}2}+1}{LT^{2}}\right)^{T}D+\frac{\zeta^{\prime\prime 2}}{2LT}.\,$ Here, $D=F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*}),{\sigma^{2}=\sum_{i\in\mathcal{S}}\alpha_{i}^{2}\sigma_{i,t}^{2}}$
@@ 6.2 FU for Balancing Local Fairness
P4: $\min_{\boldsymbol{w}}F_{-\mathcal{J}}(\boldsymbol{w}).\quad\text{s.t. }\Delta f_{i}=f_{i}(\boldsymbol{w})-f_{i}(\boldsymbol{\boldsymbol{w}}^{o})\leq\epsilon,\forall i\notin\mathcal{J}\,,$
FU Mechanism Design: To solve the above problem, we introduce Algorithm 3, adapting the saddle point optimizations as in [2, 17], using a Lagrangian multiplier $\lambda_{i}$ for each constraint: P5: $\min_{\boldsymbol{w}}\max_{\boldsymbol{\lambda}\in\mathcal{R}^{Z},\|\boldsymbol{\lambda}\|\leq\Lambda}F_{-\mathcal{J}}(\boldsymbol{w})+\boldsymbol{\lambda}^{\top}\boldsymbol{r}(\boldsymbol{w})$, where $\boldsymbol{r}(\boldsymbol{w})=\left[\Delta f_{i}-\epsilon\right]_{i\notin\mathcal{J}}$. The detailed experimental examination are provided in Section 7.
@@
This lemma can be derived from Hu et al. [17, Theorem 1]. It indicates the required unlearning rounds to reach a $\nu$-approximate saddle point of P5. It highlights that increased data heterogeneity among remaining clients and stringent fairness constraints require more unlearning rounds $T$ to balance the trade-off.
@@
Given $\epsilon=\frac{1}{\Lambda}(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}^{*})$ and assuming the existence of $\nu$-approximate saddle points of the trade-off fairness problem, then the unlearning verification metric $V\leq 2\nu=4\rho^{2}\Lambda$ and $\max_{i\notin\mathcal{J}}\Delta f_{i}\leq\epsilon$.
@@
    Data Hetegeneity: When the original FL model and optimal FU model are homogeneous, then $\epsilon=\frac{1}{\Lambda}\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{*})\right)$ as $\boldsymbol{w}^{*}$ is identical to $\boldsymbol{w}^{r*}$ under homogeneity. There is no fairness loss from unlearning but from data heterogeneity among remaining clients (as discussed in Theorem 5.11). However, under higher heterogeneity, characterized by large $F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{*})$, achieving $2\nu$ verification requires FU compromises fairness $\epsilon$.
@@
6.3 Discussion
This section demonstrates how our proposed mechanism enhances the adaptability and robustness of existing unlearning algorithms like rapid retraining [50] and knowledge distillation [49]. Specifically, introducing $\lambda$ in Wu et al. [50, Equation 3] or Wu et al. [50, Algorithm 1, Line 1] allows managing stability and unlearning effectiveness. ($\lambda=1$ prioritizes unlearning, $\lambda=0$ reverts to original FL training). Similarly, by altering the objective function with the fairness-constrained objective in P5, our proposed mechanism incorporates fairness considerations into existing unlearning algorithms.
7 Experiemnts